OpenStudy (anonymous):

(6x^4/2xy^3)^3= 36x^12/y^9 27x^9/y^9 3x^11y^9

OpenStudy (campbell_st):

well you need to recognise that the power of 3 acts on every part. You also need to know the index law for power of a power. Which requires you to multiply powers \[(x^a)^b = x^{a \times b}\] the easiest way to attempt this problem is to simplify the terms inside the brackets first \[\frac{6x^4}{2xy^3} = \frac{3x^3}{y^3}\] which saw a common factor of 2x removed now apply the power of a power law \[(\frac{3x^3}{y^3})^3=\frac{3^3 (x^3)^3}{(y^3)^3}\] which is \[\frac{27(x^3)^3}{(y^3)^3}\] now just use the power of a power rule above.